OpenStudy (anonymous):

how would you work out How many positive integers less than 100, have a remainder of 3 when divided by 7?

OpenStudy (raden):

that numbers can be represented as x = 7m + 3 just trials for m integers, get m =1,2,3,4,...,13 so, there are 13 numbers

OpenStudy (anonymous):

well I got that problem from my SAT book and it says the answer is 14 and I do not understand why

OpenStudy (raden):

if m = 14, so x = 7 * 14 + 3 = 98 + 3 = 101 that's more than 100